OpenStudy (anonymous):

y + 9 = 47 y = please help

OpenStudy (anonymous):

38

OpenStudy (anonymous):

38

OpenStudy (anonymous):

To solve for y, you want to get y by itself. In order to get y by itself, you subtracted 9 for an equation, what you do to one side, you must do to the other so y+9=47 y+9-9=47-9 y=?

OpenStudy (anonymous):

you just subtract 47 and 9

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

np :D

OpenStudy (anonymous):

(m - 5) − 16 = −12 m =

OpenStudy (anonymous):

how do i do this one ?

OpenStudy (anonymous):

The same as above, It would read as m-5-16=-12 which equals m-21=-12 this is because -5-16= -5+(-16) so you have m-21=-12 now add 21 to both sides

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ugh math is so hard for me

OpenStudy (anonymous):

y=47-9= 38

OpenStudy (anonymous):

so you should get m-21=-12 m-21+21=-12+21 m=9 You can check this answer by plugging 9 in for m into your starting equation. so (m-5)-16=-12 (9-5)-16=-12 4-16=-12 -12=-12

OpenStudy (anonymous):

-8z = -64 z =

OpenStudy (anonymous):

divide by -8

OpenStudy (anonymous):

i got a answer that idk if its right

OpenStudy (anonymous):

y + 9 = 47 y = 47-9=38 so y=38